# Dexterous Hand Manipulation via Efficient Imitation-Bootstrapped Online Reinforcement Learning

### v1

@@ III-E Online Training Algorithm
  • Lines 22-24: Apply regularization with the pretrained policy.
@@
IV EXPERIMENTS
In this section, we discuss our experiments aimed at evaluating the IBORL method. We first outline our experimental settings. We then present our experimental results and conduct an ablation analysis to investigate the effects of various components.
IV-A Experimental Settings
In our experiments, we deploy our policies on the Inspire robot, which includes a dexterous hand with six degrees of freedom, to perform various object manipulation tasks using its right arm and hand. For all tasks, the observation space consists of camera images captured from the robot’s chest and head, along with proprioceptive information, such as the end-effector pose and the angles of the robot’s fingers. The action space corresponds to the proprioceptive information found in the observation space. Specifically, the end-effector pose is represented as the relative position with respect to its initial position at the beginning of the task, comprising 3D Cartesian displacement and 3D rotational deviation using axis-angle representation.
During the policy deployment phase, we trained a binary classifier to serve as a reward detector. This classifier analyzes images to predict whether the current state successfully accomplishes the task. When the reward detector identifies the current state as successful, the episode concludes, and we manually reset the environment.
Finally, we present the evaluation tasks in Figure 3, accompanied by detailed descriptions provided below:
Cup Taking. In this task, the robot is required to grasp and lift a cup from a designated position on a table. The objective is to ensure a secure grip while smoothly elevating the cup. This task assesses the robot’s precision in handling lightweight objects and its ability to adjust grip strength.
Scanner Taking. In this task, the robot must pick up a scanner using a specified method. The challenge arises from the scanner’s irregular shape and designated grasping mode, necessitating careful planning of the grip and lifting strategy. This task evaluates the robot’s capability to handle objects with complex geometries and its adaptability in mastering the specified grasping mode.
Cube Taking. In this task, the robot is tasked with grasping and lifting a cube by its hand. Unlike the cup, the cube is smaller and heavier, thus requiring the robot to grasp the cube with its fingers. This task assesses the robot’s dexterity in manipulating its fingers.
Loopy Taking. The robot is tasked with grasping and lifting a loopy doll. This endeavor presents unique challenges due to the doll’s unconventional shape and the soft, pliable characteristics of the material.
Tasks Demos Image Inputs Reset Training Time Classifier
Cup Taking 40 Chest Camera, Head Camera Fixed 59 mins 98%
Cube Taking 36 Chest Camera, Head Camera Fixed 72 mins 96%
Scanner Taking 31 Chest Camera, Head Camera Random 41 mins 95%
Loopy Taking 36 Chest Camera, Head Camera Random 42 mins 99%
TABLE I: Task Parameters. During demo collection for both BC and RL, as well as online training, each episode’s initial end-effector pose resets uniformly at random within a fixed region
Method Cup Taking Cube Taking Scanner Taking Loopy Taking
SR (%) Cycle Time SR (%) Cycle Time SR (%) Cycle Time SR (%) Cycle Time
BC $70.0$ $122.21$ $7.5$ $129.67$ $80.0$ $204.25$ $95.0$ $116.11$
SERL 00 N/A 00 N/A 00 N/A 00 N/A
IBORL 00 N/A 00 N/A 00 N/A $100.0$ 101.8
IBORL (Ours) 87.5 68.51 90.0 110.55 90.0 147.19 100.0 $115.55$
@@ IV-A Experimental Settings
TABLE II: Experimental results comparing our method with baselines. $\dagger$ Ablated version without regularization terms. SR: Success Rate, Cycle Time: Total steps for complete the task.
IV-B Experimental Results
@@
Now, we discuss our empirical results for the real-world evaluation. For each task, we report success rate, cycle time and detailed task parameters. The training time includes all scripted motion, policy rollouts, intended stops, as well as onboard computation which is carried on a single NVIDIA RTX 4090 GPU. Unless otherwise noted, all results are based on 40 evaluation trials.
To illustrate IBORL’s superiority over existing methods, we compare IBORL with the following baselines:
  • BC: Behavior Cloning (BC) is an imitation learning method in which an agent learns to replicate an expert’s behavior by training on expert demonstrations. This approach employs supervised learning to approximate the expert’s policy.
  • SERL: Sample-Efficient Robotic Reinforcement Learning (SERL) is an open-source software framework that provides a high-quality reinforcement learning implementation specifically designed for real-world robotic learning, supporting both image observations and demonstrations.
  • IBORL: IBORL is a ablated version of IBORL, where IBORL directly applies reinforcement learning to the pretrained policy without any regularization.
@@
Results. As illustrated in Table.II, Our approach achieve competitive performance across all tasks mentioned above. Compared to BC, which serves as the initial policy for IBORL, IBORL outperforms BC by a significant margin, particularly in the tasks of Cup Taking and Cube Taking. In the Cup Taking task, IBORL not only increases the success rate from 70% to 87.5% but also reduces the cycle time required to complete the task by half. This indicates that the policy can be further optimized beyond expert performance through efficient exploration in reinforcement learning. Meanwhile, in the Cube Taking task, which requires precise finger control, BC achieves only a 7.5% success rate and an average cycle time of 129.67. This highlights that naive imitation learning struggles to master precise manipulation tasks due to compounding error issues. In contrast, IBORL effectively masters the Cube Taking task, achieving a 90% success rate and an average cycle time of 110.55 through extensive interactions with the real-world environment. This outcome demonstrates the superiority of real-world robotic reinforcement learning, as it can progressively refine its policy based on environmental feedback. Conversely, the SERL method, which focuses on real-world robotic reinforcement learning, struggles to achieve task completion due to the vast exploration space resulting from the large action space of dexterous hands and the extensive observation space derived from high-dimensional images. While IBORL benefits from imitation pretraining, successfully mastering dexterous hand manipulation skills, SERL faces significant challenges in this regard.
@@
Ablation Analysis. As illustrated in Table II, without regularization, IBORL fails to complete three tasks during real-world training. We find that reinforcement learning encounters difficulties in mastering precise and dexterous hand behaviors, such as grasping the cup, pinching the cube, and manipulating the scanner. Specifically, during the reinforcement learning process, the robotic hand tends to shake, causing the cup to be swept away and the scanner to be knocked down. Consequently, the robot is unable to complete the tasks, resulting in the absence of reward signals and leading to continued underperformance. While the Loopy Taking task, being the only one that does not require precise and complicated hand behaviors, is successfully accomplished by IBORL.
IV-C Discussions
@@
Comparison with ACT. We compare our method with Action Chunking with Transformers (ACT) to further demonstrate its superiority. ACT is a notable imitation learning algorithm that addresses compounding error issues in imitation learning by employing the transformer architecture, action chunking, and temporal ensemble techniques, resulting in significant performance improvements.
Figure 4: Comparison with ACT. In our evaluation, we conducted 40 trials per task, recording both the success rate and the average number of steps taken by the robot to achieve success in each task.
@@
As depicted in Fig.4, IBORL demonstrates competitive performance compared to the state-of-the-art imitation learning algorithm, ACT. Notably, in the Cup Taking task, IBORL accomplishes half the total steps of ACT while maintaining a comparable success rate. This indicates that reinforcement learning can surpass the limitations of expert demonstrations to derive a more optimal policy through adequate trial and error. Conversely, in the Cube Taking task, although IBORL achieves the same success rate, it is slower than ACT by an average of 20 steps. We attribute this issue to the underperforming backbone utilized in our method. In both IBORL and BC, we employ an ImageNet pre-trained ResNet-10 as the vision backbone for the policy network. In contrast, ACT utilizes the most advanced transformer architecture. When comparing BC to ACT, it is evident that although both ACT and BC share the same objective, they yield completely different performance outcomes due to the differing backbones.
@@
Regularization Analysis. To demonstrate how the regularization mechanism functions within the reinforcement learning process, we present Fig. 5, illustrating the changing trends of the balancing coefficient and the success rate. Based on this figure, we analyze the relationship between these two metrics.
@@
V CONCLUSIONS
This paper introduces IBORL, a method that leverages expert demonstrations to efficiently acquire dexterous hand manipulation skills in real-world environments. In terms of system design, IBORL establishes a framework that enhances the automation, efficiency, and compatibility of online reinforcement training. Regarding the algorithm, IBORL utilizes behavior cloning to pretrain the policy, thereby minimizing unnecessary exploration during real-world reinforcement learning. Furthermore, to tackle the challenges associated with catastrophic forgetting in reinforcement learning, we propose an adaptive regularization mechanism. Our empirical results across several real-world dexterous hand manipulation tasks demonstrate that IBORL surpasses existing methods, achieving a 1.6-fold improvement in success rate and enhanced cycle time efficiency. Additionally, through pretraining with imitation learning, IBORL successfully masters several tasks that current real-world robotic reinforcement learning methods struggle to accomplish.
Despite IBORL’s strong performance and training efficiency, we recognize that its effectiveness is significantly contingent upon the quality of the pretrained policy. In scenarios where the pretrained policy is inadequately trained, the exploration space becomes excessively vast, complicating the effective identification of sparse reward signals for reinforcement learning and thereby hindering the development of a practical policy. Consequently, our future research will concentrate on online reinforcement learning utilizing Vision-Language-Action Models in real-world applications.
ACKNOWLEDGMENT

### v2

@@ IV-D Training Pipeline
  • Lines 22-24: Apply regularization using the pretrained policy.
Tasks Demos Image Inputs Reset Training Time Classifier
Cup 40 Chest Camera, Head Camera Fixed 59 mins 98%
Cube 36 Chest Camera, Head Camera Fixed 72 mins 96%
Scanner 31 Chest Camera, Head Camera Random 41 mins 95%
Loopy 36 Chest Camera, Head Camera Random 42 mins 99%
TABLE I: Task Parameters: The initial position of the target object is randomly reset within a specified region.
V EXPERIMENTS
In this section, we discuss our experiments aimed at evaluating the DexGraspRL framework. We first outline our experimental settings. We then present our experimental results and conduct an ablation analysis to investigate the effects of various components.
V-A Experimental Settings
Hardware In our experiments, we deploy our policies on the Inspire robot, which includes a dexterous hand with 6 degrees of freedom, to perform various dexterous grasping tasks using its right arm and hand. For all tasks, the observation space consists of camera images captured from the robot’s chest and head, along with proprioceptive information, such as the end-effector pose and the angles of the robot’s fingers. The action space corresponds to the proprioceptive information found in the observation space. Specifically, the end-effector pose is represented as the relative position with respect to its initial position at the beginning of the task, comprising 3D Cartesian displacement and 3D rotational deviation using axis-angle representation.
Figure 3: Illustration of the robot used in our experiments. The robot is tasked with grasping a cup, utilizing image observations from the ZED2 and Realsense D435 cameras as inputs to control its arm and hand for task completion.
Data Collection Unlike other robotic systems, teleoperation via keyboard is impractical. We control the robotic hand using electromagnetic field (EMF) gloves [31] that track finger movements and monitor the six degrees of freedom (6-DoF) of the wrist pose based on the SLAM algorithm. This teleoperation device enables us to manipulate the robotic hand as effortlessly as we do our own, thereby facilitating task execution. During task execution, we record camera images captured from the robot’s chest and head, along with the end-effector pose and the angles of the robot’s fingers. These data serve as expert demonstrations and constitute the training data for the reward classifier after being annotated with our automatic scripts.
Task Settings We present the visualizations of the tasks in Figure 4, along with detailed descriptions provided below:
  • Cup: In this task, the robot is required to grasp and lift a cup from a designated position on a table. This task assesses the robot’s precision in handling lightweight objects and its ability to adjust grasp strength.
  • Scanner: In this task, the robot must pick up a scanner using an affordance-aware pose. This task evaluates the robot’s capability to grasp objects with complex geometries and its adaptability in mastering the affordance-aware grasping technique.
  • Cube: In this task, the robot is tasked with pinching and lifting a cube with its fingers. This task assesses the robot’s dexterity in finger control.
  • Loopy: The robot is tasked with grasping and lifting a loopy doll. This endeavor presents unique challenges due to the doll’s unconventional shape and the characteristics of the soft materials used.
Baselines To illustrate the superiority of DexGraspRL, we compare it with the following baselines:
  • BC: Behavior Cloning (BC) is a common IL method in which an agent learns to replicate an expert’s behavior by training on expert demonstrations.
  • SERL: Sample-Efficient Robotic Reinforcement Learning (SERL) is an open-source software framework that provides a high-quality RL implementation specifically designed for real-world robotic learning.
  • DexGraspRL: DexGraspRL is a ablated version of DexGraspRL, where DexGraspRL directly applies RL to the pretrained policy without any regularization.
  • ACT: Action Chunking with Transformers (ACT) is the most advanced IL method, which learns a generative model over action sequences.
@@ V-A Experimental Settings
Method Cup Cube Scanner Loopy Average
SR CT SR CT SR CT SR CT SR CT
BC $70.0$ $122.21$ $7.5$ $129.67$ $80.0$ $204.25$ $95.0$ $116.11$ $61.4$ $143.1$
SERL 0 N/A 0 N/A 0 N/A 0 N/A 0 N/A
DexGraspRL 0 N/A 0 N/A 0 N/A 100.0 101.8 $25.0$ N/A
DexGraspRL (Ours) 87.5 68.51 90.0 110.55 90.0 147.19 100.0 $115.55$ 91.9 110.45
@@
TABLE II: Experimental results comparing our method with baselines. $\dagger$ Ablated version without regularization terms. SR: Success Rate, CT: Total steps for complete the task.
V-B Experimental Results
@@
Now, we discuss our empirical results for the real-world evaluation. For each task, we report success rate, CT and detailed task parameters. The training time includes all scripted motion, policy rollouts, intended stops, as well as onboard computation which is carried on a single NVIDIA RTX 4090 GPU. Unless otherwise noted, all results are based on 40 evaluation trials.
@@
Results. As illustrated in Table.II, Our approach achieve competitive performance across all tasks mentioned above. Compared to BC, which serves as the initial policy for DexGraspRL, our approach outperforms it by a significant margin, particularly in the Cup and Cube task. In the Cup task, DexGraspRL not only increases the success rate from 70% to 87.5% but also reduces the CT required to complete the task by half. This indicates that the policy can be further optimized beyond expert performance through efficient exploration in reinforcement learning. Meanwhile, in the Cube task, which requires precise finger control, BC achieves only a 7.5% success rate and an average CT of 129.67. This highlights that naive imitation learning struggles to master precise manipulation tasks due to compounding error issues. In contrast, DexGraspRL effectively masters the Cube task, achieving a 90% success rate and an average CT of 110.55 through extensive interactions with the real-world environment. This outcome demonstrates the superiority of real-world robotic reinforcement learning, as it can progressively refine its policy based on environmental feedback. Conversely, the SERL method, which focuses on real-world robotic reinforcement learning, struggles to achieve task completion due to the vast exploration space resulting from the large action space of dexterous hands and the extensive observation space derived from high-dimensional images. While DexGraspRL benefits from BC pretraining, successfully mastering dexterous hand grasping skills, SERL faces significant challenges in this regard.
@@
Ablation Analysis. As illustrated in Table II, without regularization, DexGraspRL fails to complete three tasks during real-world training. We observe that RL encounters challenges in mastering precise and dexterous hand movements, such as grasping the cup, pinching the cube, and grabbing the scanner. Specifically, during the real-world RL training, the robotic hand tends to exhibit instability, causing the cup to be displaced and the scanner to be toppled. Consequently, the robot is unable to complete the tasks, resulting in a lack of reward signals and leading to persistent underperformance. In contrast, the Loopy task, which does not require precise and intricate hand movements, is successfully accomplished by DexGraspRL. Compare DexGraspRL to BC, we observe a rapid decline in performance following RL fine-tuning without regularization, resulting in an average degradation of 36.4% in the success rate. This highlights the importance of a regularization mechanism. Compare SERL to DexGraspRL and DexGraspRL, we observe that while SERL demonstrates impressive performance in gripper robots, it is ineffective in learning tasks without prior pretraining through IL.
V-C Discussions
@@
Comparison with ACT. We compare our method with Action Chunking with Transformers (ACT) to further demonstrate its superiority. ACT is a notable IL algorithm that addresses compounding error issues in IL by employing the transformer architecture, action chunking, and temporal ensemble techniques, resulting in significant performance improvements.
Tasks BC ACT DexGraspRL
SR CT SR CT SR CT
Cup $70.0$ $122.21$ 90.0 $102.33$ $87.5$ 68.51
Cube $7.5$ $129.67$ 90.0 91.11 90.0 $110.55$
Scanner $80.0$ $204.25$ $60.0$ 86.17 90.0 $147.19$
Loopy $95.0$ $116.11$ 100.0 85.4 100.0 $115.55$
Average $61.4$ $143.1$ $85.0$ 91.25 91.9 $110.45$
TABLE III: Comparison results with ACT.
@@
As depicted in Table.III, DexGraspRL demonstrates competitive performance compared to the state-of-the-art IL algorithm, ACT. Notably, in the Cup task, DexGraspRL accomplishes half the total steps of ACT while maintaining a comparable success rate. This indicates that RL can surpass the limitations of expert demonstrations to derive a more optimal policy through adequate trial and error. Conversely, in the Cube task, although DexGraspRL achieves the same success rate, it is slower than ACT by an average of 20 steps. We attribute this issue to the underperforming backbone utilized in our method. In both DexGraspRL and BC, we employ an ImageNet pre-trained ResNet-10 as the vision backbone for the policy network. In contrast, ACT utilizes the most advanced transformer architecture. When comparing BC to ACT, it is evident that although both ACT and BC share the same objective, they yield completely different performance outcomes due to the differing backbones.
@@
Regularization Analysis. To demonstrate how the regularization mechanism functions within the RL finetuning, we present Fig. 5, illustrating the changing trends of the balancing coefficient and the success rate. Based on this figure, we analyze the relationship between these two metrics.
@@
VI CONCLUSIONS
This paper introduces DexGraspRL, a reinforcement learning framework that directly acquires dexterous grasping skills in real-world environments. In terms of system design, DexGraspRL establishes a framework that ensures the automation, efficiency, and compatibility of online RL training. Regarding the algorithm, DexGraspRL utilizes BC to pretrain the policy, thereby mitigating the exploration challenges arising from the vast configuration space of dexterous hands. Furthermore, to address the challenges associated with catastrophic forgetting in RL fine-tuning, we propose an adaptive regularization mechanism. Our empirical results across several real-world dexterous grasping tasks demonstrate that DexGraspRL surpasses existing methods, achieving a 1.6-fold improvement in success rate and enhanced cycle time efficiency. Additionally, DexGraspRL successfully masters several tasks that current real-world robotic reinforcement learning methods struggle to accomplish.
Despite DexGraspRL’s strong performance and efficiency, we recognize that its effectiveness is significantly contingent upon the quality of the pretrained policy. In scenarios where the pretrained policy is inadequately trained, the exploration space becomes excessively vast, complicating the effective identification of sparse reward signals for reinforcement learning and thereby hindering the development of a practical policy. Therefore, our future research will concentrate on employing additional priors to enhance the sample efficiency of real-world RL training.